OpenStudy (jamel):

i need help solving 2/3b-2 > 10

jimthompson5910 (jim_thompson5910):

\[\large \frac{2}{3}b-2 > 10\] \[\large \frac{2}{3}b > 10+2\] \[\large \frac{2}{3}b > 12\] \[\large 2b > 12*3\] \[\large 2b > 36\] \[\large b > \frac{36}{2}\] \[\large b > 18\]